OpenStudy (cutebunny12):

Graph the inequality three times n minus four is less than or equal to twenty-three.

OpenStudy (anonymous):

DESMOS.COM USE THERE graphing calculator

OpenStudy (anonymous):

desmos.com

OpenStudy (cutebunny12):

ok now what

OpenStudy (anonymous):

@ROMNAY is right. Desmos Graphing Calculator is very good. It automatically graphs the coordinates you type. The only thing to watch out for is that you HAVE TO put parentheses around the coordinates for it to work.

OpenStudy (calculusxy):

\[3n - 4 \le 23\]

OpenStudy (calculusxy):

You would need to solve for n.

OpenStudy (calculusxy):

Do you know how to?

OpenStudy (cutebunny12):

yes

OpenStudy (calculusxy):

I am not sure whether you need to graph on a number line or on a coordinate graph.

OpenStudy (calculusxy):

Would you, by any chance. know?

OpenStudy (cutebunny12):

number line

OpenStudy (calculusxy):

Okay. I can help with that. Solve for n first

OpenStudy (cutebunny12):

ok

OpenStudy (calculusxy):

what did you get?

OpenStudy (cutebunny12):

hold on

OpenStudy (cutebunny12):

16

OpenStudy (calculusxy):

it's basically like solving for n in a typical equation: let's replace, for now, the less than or equal to sign with the equal sign \[3n - 4 = 23\] what would you get for n?

OpenStudy (cutebunny12):

n=1

OpenStudy (calculusxy):

are you sure? you need to isolate n so you need to get rid of 4 by doing the inverse operation of subtraction which is addition \[3 n - 4 + 4 = 23 + 4\] \[3n = 27\]

OpenStudy (cutebunny12):

i did it wrong hold on

OpenStudy (calculusxy):

Did you get it now?

OpenStudy (calculusxy):

I am going to come back after about 20 minutes. If you can handle the question yourself, then just say so in the comment section and I will know. If you can't, then only way possible is for you to be here for twenty minutes or come back after 20 minutes if you want my help. I am pretty sure that others might help you as well :)

OpenStudy (cutebunny12):

ok

OpenStudy (cutebunny12):

n=3

Vocaloid (vocaloid):

not quite we have: 3n = 27 divide both sides by 3, what do you get?